# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v1

@@ 4.3.3 Marginal Volatility Estimator
Figure 5: The figure shows how the Deep NN (Model 3) transforms the standard deviation of the rescaled returns. The x-axis shows the standard deviation of the rescaled returns, while the y-axis shows the output of the NN. The marginal distributions of the input and output are shown on the top and right sides of the figure, respectively.
@@ 4.4.1 Benchmarking Alternative Estimators
For clarity, we distinguish between methods that rely exclusively on univariate information and those that exploit the full covariance structure.
@@
Univariate benchmarks serve as low-information baselines. The Inverse-Volatility Portfolio (IVP) assigns each asset a weight proportional to the reciprocal of its estimated variances, thereby equalising the contribution of each asset to total ex-ante volatility; the idea was popularised by Maillard, Roncalli and Teiletche [61]. The capitalisation-weighted portfolio (MCAP) coincides with the market proxy used in empirical asset-pricing work and therefore absorbs only the cross-sectional information embedded in free-float market values.
Multivariate benchmarks differ only in the way the $n\times n$ correlation matrix $\widehat{\bm{C}}$ is obtained. After the correlation matrix is estimated, we always project on the univariate volatilities $\widehat{\bm{\Sigma}}=\widehat{\bm{D}}\widehat{\bm{C}}\widehat{\bm{D}}$ and we apply the QP optimization. Filtering the correlation rather than the covariance is a necessary step to put RIEs in their best performance setup.
@@
The simplest method is the sample MLE uses the unfiltered sample correlation matrix computed over the look-back window. The Power Mapping (PM), introduced by Guhr and Kälber [62], suppresses estimation noise by raising every off-diagonal element of the de-meaned correlation matrix to a power $\gamma>1$ and re-normalising; following their empirical recommendations, we set $\gamma=1.2$. For the LS, we use the standard implementation of Sklearn [63] where the parameter is set up according to the original work of Ledoit and Wolf [10]. For QIS, we used a Python implementation released in the GitHub repository [64] that we adapted to improve the computational performance. Finally, AO is calibrated using the universe of the stocks as defined in Sec. 4.1 within 1990-2000. Then the eigenvalues are kept constant for the whole validation period, 2000-2024, independently of the stock in the basket.
With the exception of the MCAP, which is rebalanced monthly, all the other portfolios are rebalanced at the same weekly frequency as the NN strategy, and they are compared on the same set of stocks.
@@ 4.4.2 Frictionless Interpolation Backtesting
To assess in a statistically robust way the performances, we performed a battery of $1,000$ independent backtests. Specifically, every simulation consists of approximately $5$ years of data, starting from a random day $t$ within [2000-01-01, 2019-12-31], and performing $250$ portfolio rebalancing every $5$ trading days. Our model will use the last available calibration obtained the year before the trading day so that no data leakage occurs. The list of assets is randomly sampled from the investable universe (see Sec. 4.1) on the first day of the simulation, and it is kept fixed if possible for the entire backtest. If some of the assets are excluded from the investable universe during the backtest, they are removed from the portfolio, and another random asset is added to the portfolio so that the number of assets remains constant.
Method SR Volatity Mean Return $n_{\text{eff}}$ Turnover Gross Leverage
NN 1.011 0.109 0.110 15.2 1.125 3.07
QIS 0.942 0.111 0.105 10.6 0.440 3.81
AO 0.907 0.121 0.110 28.6 0.163 2.13
PM 0.902 0.114 0.102 10.3 0.534 4.00
LS 0.893 0.115 0.103 7.2 0.708 4.74
MLE 0.868 0.117 0.102 6.5 0.785 5.01
Table 1: Unconstrained long-short portfolio backtests. The table shows the average performances of $1,000$ independent simulations of the experiments with 300 stocks within [2000-01-01, 2019-12-31] performing $250$ portfolio rebalancing every $5$ trading days, without transaction costs. Values are sorted according to the Sharpe Ratio (SR), and bolds represent statistically significant top rank according to a percentile bootstrap with p-value smaller than $10^{-6}$. Portfolio volatility, mean returns and SR are annualized, $n_{\text{eff}}$ is computed by using the inverse Herfindahl index [65].
Method SR Volatility Mean Return $n_{\text{eff}}$ Turnover
NN 0.792 0.135 0.107 16.5 0.469
AO 0.740 0.146 0.108 33.2 0.098
LS 0.723 0.142 0.103 13.6 0.142
PM 0.721 0.143 0.103 17.6 0.131
MLE 0.721 0.142 0.102 13.3 0.143
QIS 0.719 0.142 0.102 13.5 0.138
Table 2: Constrained long-only portfolio backtests. The table shows the average performances of $1,000$ independent simulations of the experiments with 300 stocks within [2000-01-01, 2019-12-31] performing $250$ portfolio rebalancing every $5$ trading days, without transaction costs. Values are sorted according to SR, and bolds represent statistically significant top rank according to a percentile bootstrap with p-value smaller than $10^{-6}$. Portfolio volatility, mean returns and SR are annualized, $n_{\text{eff}}$ is computed by using the inverse Herfindahl index [65].
In Tabs. 1 and 2 we report the average performances of the models tested on the same set of assets. The unconstrained (long–short) results in Tab. 1 show that the NN achieves the highest SR (1.01) with an annualized volatility of $10.9\%$ and a mean return of $11.0\%$. The QIS estimator attains the second-best SR ($0.94$) but does so with a notably high gross leverage of $3.81$, compared with the NN leverage of 3.07. The AO method ranks third in SR ($0.91$) yet operates with the lowest leverage ($2.13$) and minimal turnover ($16.3\%$). These performance gains for the NN and QIS come at the cost of trading intensity. The network rebalances more than its entire notional every five trading days (turnover $112.5\%$), and QIS rebalances $44.0\%$ of its portfolio at each rebalance. Such high turnover and leverage would incur significant financing, borrowing, and market-impact costs not captured in this frictionless experiment.
Once short-selling is forbidden (Tab. 2), most of the models perform similarly. This is a known behavior of covariance cleaning for portfolio optimization [26, 44]. The annualized volatility is around $14.2\%$for all the models, including the MLE covariance estimator, which is the non-cleaned covariance matrix. The AO estimator performs slightly worse, reaching $14.6\%$ annualized volatility. The NN instead reaches $13.5\%$ annualized volatility, which is the lowest among all the models by a substantial margin. In terms of the SR, the NN reaches a $0.79$, which is the highest among all the models, while the AO estimator reaches the second-best SR of $0.74$, while all the other methods range around $0.72$. We want to stress that checking the SR is very important when measuring the performance of a portfolio in a large, uncontrolled setting. In fact, it might be possible that the data cleaning included stocks that must be excluded, for example, stocks that are about to be delisted or merged or stocks that are not liquid enough. In those cases the asset might have a very low volatility and a very low return. A portfolio that includes those assets will have a very low volatility, but it will also have a very low return, leading to a very low SR. The objective of a GMV portfolio is, in fact, to minimize the volatility of investing in risky assets.
@@
The other two columns in Tab. 2 report the average number of effective assets in the portfolio. This metric is dominated by the AO estimator, which has an average of $33.2$ effective assets, while the others range between $13.3$ and $17.6$, which are less than half of the number of assets in the portfolio.
Finally, the average turnover reported in the last column measures the fraction of assets that are changed in the portfolio at every rebalancing. The AO estimator has the lowest turnover ($9\%$) while the NN has a turnover of $47\%$, which is the highest among all the models. This is due to the fact that the NN is trained to be reactive and to focus on the most recent past, while the other models are more stable and less sensitive to recent changes. This behavior is a double-edged sword; on one side, having a more reactive portfolio allows one to adapt to the market conditions, but on the other side, it leads to a higher turnover and, therefore, higher transaction costs. For this reason, in the next section, we will evaluate the performance of the NN in a more realistic setting.
@@ 4.4.3 High Realistic Extrapolation Backtesting
At any date $t$, the model keeps a position vector $\mathbf{s}_{t}=\left\{s_{t,i}\right\}_{i=1}^{n}$ expressed in shares and a free cash $c_{t}$ scalar. Cash dividends are credited automatically and increase $c_{t}$. If free cash turns negative after trading, the account is deemed in debit. Interest accrues daily on the debit balance according to the IBKR tiered schedule [66]. Conversely, when $c_{t}$ is positive, the broker pays credit interest on the portion above successive notional caps according to the corresponding credit-tier spreads; however, since the first 100,000\$ tranche is unpaid, we are not modeling that. To decide the daily rate a 360-day convention is used, so the daily debit rate is the FED funds rate $\phi^{(r)}_{t}$ [67] plus a spread $\delta_{v}$, i.e., $\phi_{t}=\phi^{(r)}_{t}+\delta_{v}$. While the spread $\delta_{v}$ can be found on the IBKR website [68], the reference rate $\phi^{(r)}_{t}$ is a proprietary algorithm of the broker, and it is not available to the public. Nevertheless, we think that the FED funds rate could be a good proxy. It is worth noting that on the tiered schedule, the spread depends on the traded volume of the account in the current month.
@@
The difference between the target number of shares $\bm{s}_{t}$ and the previous-day position $\bm{s}_{t-1}$, corrected for eventual splits, is used as an end-of-day market order request $\bm{m}_{t}=\bm{s}_{t}-\bm{s}_{t-1}$. The trades are executed with the actual close price of the day $\bm{p}_{t}$, without accounting for market impact.
@@
The simulator returns, for every calendar date, the full path of $\text{NLV}_{t}$, cash, market values, and margin statistics, thus reflecting the realistic interaction between turnover, leverage due to slippage, and transaction costs. The simulation starts with a cash balance of 1,$0$, which is a reasonable amount for a small hedge fund portfolio, and it is sufficiently low to avoid the need to account for market impact on the most liquid US stocks. Our portfolio is rebalanced every $5$ trading days, and the investable asset universe is taken from the procedure described in Sec. 4.1, with the only difference that we select the top $n=1,000$ for marketcap, defined from the previous day’s close price. In this case, we are attempting to extrapolate the NN to a larger number of assets, given that the calibration was done on a maximum of $350$ assets. Furthermore, $q=n/\Delta t_{\text{in}}\approx 0.83$, which is substantially larger than the training range of $q\in[0.04,0.29]$.
SR Mean Return Volatility Max Drawdown Turnover $n_{\text{eff}}$
NN 1.046 0.122 0.117 -0.350 0.589 33.4
AO 0.951 0.118 0.124 -0.374 0.179 45.4
PM 0.880 0.109 0.124 -0.387 0.224 23.9
QIS 0.859 0.107 0.125 -0.388 0.230 19.0
MLE 0.858 0.107 0.125 -0.385 0.248 17.9
IVP 0.678 0.123 0.181 -0.504 0.058 728.0
MCAP 0.483 0.094 0.195 -0.542 0.065 133.1
@@
Table 3: Performance of the long-only portfolio with $n=1000$ top capitalization stocks, rebalanced every 5 trading days, from January 2000 to December 2024. The simulations use a realistic trading simulator that accounts for transaction costs and slippage. SR, returns, and volatility are annualized. The turnover metric is computed without accounting for price drift and serves only as a reference to the model’s stability, $n_{\text{eff}}$ is computed by using the inverse Herfindahl index [65].
@@
In Tab. 3, we can observe the performance metrics of the portfolio across different strategies. The NN strategy outperforms the others in terms of SR, volatility, and maximum drawdown, and reaches the second-best mean return. The second-best SR is achieved by the AO estimator. All the other methods, including the MLE covariance estimator, have similar performances in terms of SR and volatility. The univariate methods, such as IVP and MCAP, perform significantly worse than the multivariate methods, with an SR of $0.678$ and $0.483$, respectively. In fact, for the IVP strategy, which allocates weights by inverting individual variances in a univariate fashion, the diversification indicator $n_{\text{eff}}$ reaches 728 out of 1 000, reflecting an almost equally weighted allocation. As detailed in Sec. 4.1, stocks with exceptionally low IS volatility were excluded, thereby preserving the multivariate nature of the optimization problem.
@@
In the top left panel of Fig. 6, we show the NLV of the portfolio as a function of time. The NLV is computed as in Eq. (21) and it is shown in log scale. From the figure, we can observe that the better performance of the NN in terms of returns is even higher after 2012. This is very important since many NN-based approaches tend to outperform only in the far past, typically before the huge surge of artificial intelligence [69], but we observe the exact opposite. In the plot, we just focus on four models, the NN, AO, PM, and MLE, which are the ones that are most different from each other according to the performance metrics in Tab. 3. In the top right panel of Fig. 6, we show the annualized volatility of the NN and AO portfolios as a function of time over a yearly rolling window. The NN portfolio has a systematically lower volatility than AO across almost the entire period, with the only exception of 2020, which is the year of the COVID-19 pandemic. In the bottom panel of Fig. 6, we show the Maximum Drawdown (MDD) of the portfolio by year. Also, this plot confirms that the NN portfolio has a systematically lower MDD than AO, with only a few exceptions.
@@
Figure 6: Long-only portfolio performances of the top 1000 most capitalized stocks in the universe backtested with the high realistic simulator. Top left: NLV of the portfolio as a function of time. Top right: Annualized volatility of the portfolio as a function of time over a yearly rolling window. Bottom: MDD of the portfolio by year.

### v2

@@ 4.3.3 Marginal Volatility Estimator
Figure 5: The figure shows how the MLP (Model 3) transforms the standard deviation of the lag-transformed returns. The x-axis shows the standard deviation of the lag-transformed returns, while the y-axis shows the output of the NN. The marginal distributions of the input and output are shown on the top and right sides of the figure, respectively.
@@ 4.4.1 Benchmarking Alternative Estimators
We compare two broad classes of portfolio estimators, (i) univariate benchmarks and (ii) multivariate covariance-cleaning methods, across two complementary experimental designs introduced in Secs.  4.4.2 and 4.4.3. In Sec.  4.4.2 conduct a bootstrap-style backtest without transaction costs to enable frictionless, synthetic comparisons over varying universe sizes. In Sec. 4.4.3, we perform a single, realistic simulation on the top-capitalisation universe of 1,000 stocks, incorporating trading frictions, liquidity constraints, and slippage. Univariate benchmarks rely solely on individual asset variances or market–weight information. Multivariate methods instead apply covariance-cleaning procedures to the estimated correlation matrix prior to optimization. Although univariate strategies are pervasive in index construction and passive instruments, their ex post performance is predictably inferior; accordingly, we include them only in our detailed analysis of the realistic backtest of Sec. 4.4.3.
@@
Univariate benchmarks serve as low-information baselines. The Equal-Risk Budget (ERB) assigns each asset a weight proportional to the reciprocal of its estimated variance, thereby equalising the contribution of each asset to total ex-ante volatility [72]. The Market-Capitalization Weighted (MCW) portfolio coincides with the market proxy used in empirical asset-pricing work and therefore absorbs only the cross-sectional information embedded in free-float market values [73].
@@
Multivariate benchmarks differ only in the way the $n\times n$ correlation matrix $\widehat{\bm{C}}=\bm{C}_{\bullet}$ is obtained. After the correlation matrix is estimated, we always project on the univariate volatilities $\widehat{\bm{\Sigma}}=\bm{D}_{\textrm{MLE}}\bm{C}_{\bullet}\bm{D}_{\textrm{MLE}}$ and we apply the QP optimization. Filtering the correlation rather than the covariance is a necessary step to put RIEs in their best performance setup. The simplest method is the sample MLE uses the unfiltered sample correlation matrix computed over the look-back window. The Power Mapping (PM), introduced by Guhr and Kälber [74], suppresses estimation noise by raising every off-diagonal element of the de-meaned correlation matrix to a power $\gamma>1$ and re-normalising. For the LS, we use the standard implementation of Sklearn [75] where the parameter is set up according to the original work of Ledoit and Wolf [6]. For QIS, we used a Python implementation released in the GitHub repository [76] that we adapted to improve the computational performance. Finally, AO is calibrated using the universe of stocks as defined in Sec. 4.1 within 1990-2000. Then the eigenvalues are kept constant for the whole validation period, 2000-2024, independently of the stock in the basket.
We exclude dynamical models, most notably the DCC estimator and its myriad extensions, from our benchmark analysis. Although a vast plethora of DCC variants has emerged in the literature (see Ref. [77]), a thorough comparison against all these methods would require a dedicated study: the differing model formulations and hyper-parameter protocols alone would demand a tailored experimental framework, and the well-documented computational overhead and convergence instabilities of DCC algorithms further compound the effort. More fundamentally, none of the estimators considered in this work, including our end-to-end NN architecture, are designed to produce genuinely time-evolving covariance matrices; they all operate within a static eigenvalue-filtering and shrinkage paradigm. Incorporating DCC-based methods would therefore conflict with the conceptual scope of our paper. A coherent evaluation would entail extending our network with modules to map sequences of covariance matrices and then re-benchmarking it against the full spectrum of DCC approaches, a task beyond the present work. Moreover, Ref. [27] already shows that a standard DCC implementation underperforms the AO estimator on the same universe of the top 1,000 U.S. large-cap equities, with calibration window $\Delta t_{\textrm{in}}=1200$, yielding consistently lower Sharpe ratios for DCC in realistic settings. This evidence further justifies our focus on static cleaning and shrinkage procedures.
@@ 4.4.2 Frictionless Interpolation Backtesting
To assess the performance in a statistically robust way, we performed a battery of 1,0001,0001 , 000 independent backtests. Specifically, every simulation consists of approximately 555 years of data, starting from a random day $t$ within [2000-01-01, 2019-12-31], and performing 250250250 portfolio rebalancing every 555 trading days. Our model will use the last available calibration obtained the year before the trading day so that no data leakage occurs. The list of assets is randomly sampled from the investable universe (see Sec. 4.1) on the first day of the simulation, and it is kept fixed if possible for the entire backtest. If some of the assets are excluded from the investable universe during the backtest, they are removed from the portfolio, and another random asset is added to the portfolio so that the number of assets remains constant.
Method $\mathcal{L}(\bm{w},\bm{\Sigma}_{\mathrm{out}})$ Volatility Mean Return Sharpe Ratio $\mathbf{n_{\mathrm{eff}}}$ Turnover Leverage
Unconstrained (Long –Short)
NN 0.0149 10.9% 11.0% 1.011 15.2 112.0% 3.07
QIS 0.0151 11.1% 10.5% 0.942 10.6 44.0% 3.81
AO 0.0184 12.1% 11.0% 0.907 28.6 16.0% 2.13
PM 0.0158 11.4% 10.2% 0.902 10.3 53.0% 4.00
LS 0.0163 11.5% 10.3% 0.893 7.2 71.0% 4.73
MLE 0.0168 11.7% 10.2% 0.868 6.5 78.0% 5.02
Constrained (Long –Only)
NN 0.0234 13.5% 10.7% 0.792 16.5 47.0% 1.00
AO 0.0274 14.6% 10.8% 0.740 33.2 10.0% 1.00
LS 0.0257 14.2% 10.3% 0.723 13.6 14.0% 1.00
PM 0.0259 14.3% 10.3% 0.721 17.6 13.0% 1.00
MLE 0.0257 14.2% 10.2% 0.721 13.3 14.0% 1.00
QIS 0.0258 14.2% 10.2% 0.719 13.5 14.0% 1.00
Table 1: Performance comparison of frictionless portfolio backtests for both unconstrained (long-short) and constrained (long-only) settings. The table reports averages over 1,0001{,}0001 , 000 independent simulations of 250250250 rebalancings (every 5 trading days) on portfolios of $n=300$ stocks randomly sampled during [2000-01-01,2019-12-31], without transaction costs. Values are sorted by Sharpe ratio, and bold entries indicate the statistically significant top rank (percentile bootstrap, $p<10^{-6}$). Portfolio volatility, mean return, and Sharpe ratio are annualized; $n_{\mathrm{eff}}$ is computed via the inverse Herfindahl index [78]. The loss function $\mathcal{L}(\bm{w},\bm{\Sigma}_{\mathrm{out}})$ is defined in Eq.(13) and refers to the short-term variance ($\Delta t_{\textrm{out}}=5$ days).
[…]
@@
The other two columns in Tab. 1 report the average number of effective assets in the portfolio. This metric is dominated by the AO estimator, which has an average of 33.233.233.2 effective assets, while the others range between 13.313.313.3 and 17.617.617.6, which are less than half of the number of assets in the portfolio.
[…]
@@ 4.4.3 High Realistic Extrapolation Backtesting
At any date $t$, the model keeps a position vector $\mathbf{s}_{t}=\left\{s_{t,i}\right\}_{i=1}^{n}$ expressed in shares and a free cash $c_{t}$ scalar. Cash dividends are credited automatically and increase $c_{t}$. If free cash turns negative after trading, the account is deemed in debit. Interest accrues daily on the debit balance according to the IBKR tiered schedule [79]. Conversely, when $c_{t}$ is positive, the broker pays credit interest on the portion above successive notional caps according to the corresponding credit-tier spreads; however, since the first \$100,000 tranche is unpaid, we do not model that. To determine the daily rate, a 360-day convention is used. IBKR defines its USD reference rate $\phi^{(r)}_{t}$ as the Fed Funds Effective Rate i.e., the volume-weighted average of interbank transactions processed through the Federal Reserve [80], so we set $\phi_{t}=\phi^{(r)}_{t}+\delta_{v}$ where $\delta_{v}$ is the volume-dependent spread published by IBKR [81].
@@
The simulator returns, for every calendar date, the full path of $\text{NLV}_{t}$, cash, market values, and margin statistics, thus reflecting the realistic interaction between turnover, leverage due to slippage, and transaction costs. The simulation starts with a cash balance of 1,$0$, which is a reasonable amount for a small hedge fund portfolio, and it is sufficiently low to avoid the need to account for market impact on the most liquid US stocks [82]. Our portfolio is rebalanced every 555 trading days, and the investable asset universe is taken from the procedure described in Sec. 4.1, with the only difference that we select the top $n=1,000$ for marketcap, defined from the previous day’s close price. In this case, we are attempting to extrapolate the NN to a larger number of assets, given that the calibration was done on a maximum of 350350350 assets. Furthermore, $q=n/\Delta t_{\text{in}}\approx 0.83$, which is substantially larger than the training range of $q\in[0.04,0.29]$.
Method Short Var. Sharpe Ratio Sortino Ratio Mean Return Volatility Turnover $\mathbf{n_{\mathrm{eff}}}$
NN 9.0% 1.058 1.268 12.6% 11.9% 57.0% 46.0
AO 9.7% 0.942 1.155 11.7% 12.5% 18.0% 47.8
PM 9.8% 0.870 1.077 10.8% 12.5% 23.0% 25.7
QIS 9.8% 0.848 1.053 10.6% 12.6% 23.0% 20.1
MLE 9.8% 0.847 1.052 10.6% 12.5% 25.0% 18.8
ERB 20.2% 0.533 0.710 13.0% 24.5% 6.0% 580.2
MCW 15.7% 0.484 0.615 9.4% 19.5% 7.0% 146.6
@@
Table 2: Performance of the long-only portfolio with $n=1000$ top capitalization stocks, rebalanced every 5 trading days, from January 2000 to December 2024. The simulations use a realistic trading simulator that accounts for transaction costs and slippage. Ratios, returns, and volatility are annualized. The turnover metric is computed without accounting for price drift and serves only as a reference to the model’s stability, $n_{\text{eff}}$ is computed by using the inverse Herfindahl index [78]. The first column indicates the average short-term portfolio variance ($\Delta t_{\textrm{out}}=5$) consistently with the NN loss function (Eq. (13)). Bold values indicate the best-ranked model.
@@
In Tab. 2, we can observe the performance metrics of the portfolio across different strategies. Notably, when considering the realized variance over the five-day OOS window (i.e., the first column of the table), our NN model again achieves the lowest variance, a result that is fully consistent with its calibration objective. The NN strategy outperforms the others in terms of Sharpe ratio, Sortino ratio, volatility, and reaches the second-best mean return. The second-best Sharpe ratio is achieved by the AO estimator. All the other methods, including the MLE covariance estimator, have similar performances in terms of Sharpe ratio, Sortino ratio, and volatility. The univariate methods, such as ERB and MCW, perform significantly worse than the multivariate methods, with a Sharpe ratio of 0.6780.6780.678 and 0.4830.4830.483, respectively. In fact, for the ERB strategy, which allocates weights by inverting individual variances in a univariate fashion, the diversification indicator $n_{\text{eff}}$ reaches 580 out of 1,000, reflecting a relatively uniform allocation. As detailed in Sec. 4.1, stocks with exceptionally low past volatility were excluded, thereby preserving the multivariate nature of the optimization problem.
@@
In the top left panel of Fig. 6, we show the NLV of the portfolio as a function of time. The NLV is computed as in Eq. (22) and it is shown in log scale. From the figure, we can observe that the better performance of the NN in terms of returns is even higher after 2012. This is very important since many NN-based approaches tend to outperform only in the far past, typically before the huge surge of artificial intelligence [83], but we observe the exact opposite. In the plot, we just focus on four models, the NN, AO, PM, and MLE, which are the ones that are most different from each other according to the performance metrics in Tab. 2. In the top right panel of Fig. 6, we show the annualized volatility of the NN and AO portfolios as a function of time over a yearly rolling window. The NN portfolio has a systematically lower volatility than AO across almost the entire period, with the only exception of 2020, which is the year of the COVID-19 pandemic. In the bottom panel of Fig. 6, we show the Maximum Drawdown (MDD) of the portfolio by year. Also, this plot confirms that the NN portfolio has a systematically lower or equal MDD than AO, with only a few exceptions.
@@
Figure 6: Long-only portfolio performances of the top 1,000 most capitalized stocks in the universe backtested with the high realistic simulator. Top left: NLV of the portfolio as a function of time. Top right: Annualized volatility of the portfolio as a function of time over a yearly rolling window. Bottom: MDD of the portfolio by year.